OpenStudy (isabel☺):

A stadium with a field, one side of which is 100 m., contains 20,000 hectares. How many squares meters are there in a similar field whose corresponding side is 75 m?

OpenStudy (isabel☺):

@amistre64 @Vincent-Lyon.Fr

OpenStudy (arbershabani97):

100m 20000 hectares 75m x hectares _____________________ 100x=20000*75 x=20000*75/100

OpenStudy (isabel☺):

i need to convert hectares to meters right? @arbershabani97

OpenStudy (arbershabani97):

you don't have to because x=20000hec * 75m / 100m so meters cancel out and just hectares are left

OpenStudy (isabel☺):

wow right :D thanks thanks a lot!!